# Sharp one-point estimates and Minkowski content for the scaling limit of three-dimensional loop-erased random walk

### v1

@@ 1 Introduction
In considering the scaling limit of discrete random curves, it is natural to rescale the space by the inverse of the lattice mesh, parametrize the discrete curve in a way such that each lattice edge is traversed in the same amount of time and rescale the time in a way such that a smooth bounded domain is traversed in time of order 1. If the limit of the rescaled and parametrized random curve exists, we refer to it as the scaling limit in natural parametrization. Recently, the last two authors proved in [19, 20] that when rescaled properly, three-dimensional LERW converges as a parametrized curve to its scaling limit in natural parametrization.
With these results in mind, it is natural to ask if there is a more direct way to obtain the natural parametrization of the scaling limt. In the 2D case, Lawler and Viklund showed in [17, 18] that the scaling limit of planar LERW in natural parametrization is given through the Minkowski content of ${\rm SLE}_{2}$, whose existence was confirmed in [14]. It has also been shown (see [16, 2]) that the natural parametrization of ${\rm SLE}_{k}$ (which satisfies a conformal Markov property) is unique up to multiplicative constants. However, conceivably it is difficult to establish such axiomatic characterization of natural parametrization due to lack of the rich conformal symmetry in 2D.
In this work we are going show that the Minkowski content of the scaling limit of 3D LERW exists and is equivalent to the limiting occupation measure obtained by the authors in [20] from which the natural parametrization of the scaling limit is induced. In the course of proof, we also obtain sharp asymptotics of one-point function in any scale, and sharp estimates on the ball-hitting probabilities of LERW and its scaling limit, which is a big improvement of earlier results of the authors in [19].
@@ 1.1 Main results
We now turn to the main results of this paper. Before stating them precisely, we need to introduce some notation first. Let $\mathbb{D}=\{x\in\mathbb{R}^{3}\ |\ |x|<1\}$ be the open unit ball. Pick some $m\in\mathbb{R}^{+}$ and let $\mathbb{D}_{m}\coloneqq m^{-1}\mathbb{Z}^{3}\cap\mathbb{D}$ be the discretized unit ball on the rescaled integer lattice $m^{-1}\mathbb{Z}^{3}$.
Let $S_{m}$ be the simple random walk on $m^{-1}\mathbb{Z}^{3}$ started from 00, stopped at the first exit of $\mathbb{D}_{m}$ and let $\eta_{m}\coloneqq{\rm LE}(S_{m})$ be its loop-erasure which we will refer to as the LERW on $\mathbb{D}_{m}$. (We define the loop-erasure of a path below, in Definition 2.1.) Write ${\cal K}_{m}$ for the trace of $\eta_{m}$ as a random compact subset of $\mathbb{R}^{3}$ and let ${\cal K}$ stand for the weak limit of ${\cal K}_{2^{n}}$, as $n\in{\mathbb{Z}^{+}}$ tends to infinity, whose existence was confirmed in [10]. Let
@@
be the reference scaling factor. Here $\hat{x}_{m}$ stands for one of the nearest point from $\hat{x}$ among $m^{-1}\mathbb{Z}^{3}$. From up-to-constant one-point estimates, we have $f_{m}\asymp m^{\beta}$ (see Subsection 2.6 below). Let
$\mu_{m}=(f_{m})^{-1}\sum_{x\in\mathbb{D}_{m}\cap\eta_{m}}\delta_{x}.$ (1.3)
@@
It was proved in [20, Theorem 1.1] that as $m\in{\mathbb{R}^{+}}$ tends to infinity,
@@
for some random measure $\mu$ supported on ${\cal K}$ which is measurable with respect to ${\cal K}$. We call $\mu$ the limiting occupation measure.
@@
From now on we regard the LERW $\eta_{m}$ as a parametrized curve where each edge of $m^{-1}\mathbb{Z}^{3}$ is traversed in time $(f_{m})^{-1}$ with a constant speed. Let $\eta$ be the random continuous curve obtained through parametrizing ${\cal K}$ by the limiting occupation measure $\mu$ as in [20, Subsection 9.1]. It was also proved in [20, Theorem 1.3] that $\eta_{m}$ converges to $\eta$ in natural parametrization. In other words, as $m\in{\mathbb{R}^{+}}$ tends to infinity,
@@
We now turn to Minkowski content. Roughly speaking, the Minkowski content of a set to be measured is defined through the limit of the renormalized volume of the enlargement of the set. Although not a measure per se, it is a natural way to define “fractal” measures on fractal sets. Given a Borel set $G\subset{\mathbb{R}^{3}}$ and $\delta\in(0,3]$, we write ${\rm Cont}_{\delta}(G)$ for its $\delta$-dimensional Minkowski content (provided it exists); see Subsection 2.3 for its precise definition. Let ${\cal D}^{o}$ be the collection of “nice” dyadic boxes defined in (2.12).
@@
There exists some universal $c_{0}>0$ such that for any dyadic box $V\in{\cal D}^{o}$, ${\rm Cont}_{\beta}({\cal K}\cap V)$ exists and equals to $c_{0}\mu(V)$ almost surely.
By a standard measure-theoretical procedure, see Proposition 2.2 below, we can generate a random non-atomic Borel measure $\nu$ out of the Minkowski content of ${\cal K}$.
@@
The Borel measure $\nu$ is supported on ${\cal K}$ and defines an intrinsic parametrization. Given $x\in{\cal K}$, denote by ${\cal K}_{x}\subset{\cal K}$ the trace of the curve between 00 and $x$. For each $t\in[0,\nu({\cal K})]$, there uniquely exists $x_{t}\in{\cal K}$ such that $t=\nu({\cal K}_{x_{t}})$, and then we define
@@
The relevance of Corollary 1.2 is that the parametrization of $\eta$, which was obtained via the scaling limit of the occupation measure, is equal to an intrinsic parametrization of ${\cal K}$, up to a universal constant factor. Moreover, since the measure $\nu$ is scaling-invariant and rotation-invariant, then the limit curve $\eta$ is scaling-invariant and rotation-invariant as well (see Corollary 6.5).
In the course of the proof of Theorem 1.1, we also obtain accurate asymptotics for any scale of the probability that LERW passes a certain point or a certain ball which we shall both refer to loosely as “one-point functions”. We present them here since both of them are of independent interest.
@@
We now turn to the continuum. Elaborating on Theorem 1.3 and various coupling techniques developed so far, we are able to show the following asymptotics of the probability that the scaling limit ${\cal K}$ hits a ball.
@@
There exist $c_{2},c,\delta>0$ such that for any $x\in\mathbb{D}\setminus\{0\}$, for $r<\min\{|x|,1-|x|\}/2$,
$P\Big{(}{\cal K}\cap B(x,r)\neq\emptyset\Big{)}=c_{2}g(x)r^{3-\beta}\Big{[}1+O\big{(}d_{x}^{-c}r^{\delta}\big{)}\Big{]}.$ (1.9)
@@
Note that $g(x)$ here corresponds to $c(x)$ in [19, Theorem 1.1] (which we rephrase in Proposition 2.11) where only dyadic scales were considered. Hence, $g(x)$ satisfies the following asymptotics (see [19, Equations (1.5), (1.6)]):
@@
Moreover, there exists $c_{1}>0$ such that for any $(z,w)\in\Omega$ and $r<d_{z,w}/2$
$P\left({\cal K}\cap B(z,r)\neq\emptyset,{\cal K}\cap B(w,r)\neq\emptyset\right)=c_{1}g(z,w)r^{2(3-\beta)}\Big{[}1+O(d_{z,w}^{-c}r^{\delta})\Big{]},$ (1.15)
@@
Theorem 1.1 and Theorem 1.3 enhance [20, Theorem 1.3]. Specifically, the refined one-point estimates allow us to substitute the reference scaling factor (1.2) with an explicit scaling. We write
$\overline{\mu}_{m}=m^{-\beta}\sum_{x\in\mathbb{D}_{m}\cap\eta_{m}}\delta_{x}$ (1.16)
and introduce $\overline{\eta}_{m}$ as the LERW on $\mathbb{D}_{m}$ parametrized by $\overline{\mu}_{m}$. We now present the convergence of 3D LERW in natural parametrization in its full strength.
@@
As $m\to\infty$, $\eta_{m}\xrightarrow{\rm w}\eta$ in the topology generated by the $\rho$-metric.
@@
Let $\mathcal{C}$ be the collection of parametrized curves $\lambda:[0,\infty)\to\mathbb{R}^{3}$ with $\lambda(0)=0$ and $\lim_{t\to\infty}{\rm diam}(\lambda[0,t])=\infty$. Consider the metric on $\mathcal{C}$ given by
@@
Let $S_{m}$ be a simple random walk on $m^{-1}\mathbb{Z}^{3}$ started from 00. The path
@@
defines the infinite loop-erased random walk (ILERW) on $m^{-1}\mathbb{Z}^{3}$. Since the simple random walk is transient in $\mathbb{Z}^{3}$ almost surely, then $\eta^{\infty}_{m}$ is well-defined. We also write $\eta^{\infty}_{m}$ for the curve defined by linear interpolation on the path $\eta^{\infty}_{m}$, and consider a parametrization $\overline{\eta}^{\infty}_{m}$ defined by
@@
Using the same techniques employed in proving Theorem 1.3, one can show the existence of the one-point function associated to $\eta^{\infty}$, the scaling limit of the ILERW. Specifically, there exist a function $g_{\infty}:\mathbb{D}\setminus\{0\}\rightarrow\mathbb{R}^{+}$ and a constant $c_{1}>0$ such that for any $x\in\mathbb{D}\setminus\{0\}$ and $r<d_{x}/2$,
$P\left(\eta^{\infty}\cap B(x,r)\neq\emptyset\right)=c_{1}g_{\infty}(x)r^{3-\beta}\Big{[}1+o(1)_{s}\Big{]}.$
Furthermore, the scaling invariance of $\eta^{\infty}$ implies a explicit form for $g_{\infty}$: there exists a constant $c_{2}>0$ such that $g_{\infty}(x)=c_{2}|x|^{-(3-\beta)}$.
@@
The existence of a scaling limit for the uniform spanning tree on ${\mathbb{Z}}^{3}$ has been established in [1]. The paper relies on the scaling limit of the loop-erased random walk in [20] and the one-point function estimates for the loop-erased random walk in [19]. Building upon the results from these papers, [1] shows the existence of sub-sequential scaling limits, convergence over a dyadic scaling, and properties of the limit tree.
As noted in [1, Remark 1.2], the restrictions of [20] and [19] to a dyadic scaling impose the same condition over the results in [1]. With the introduction of Theorems 1.31.7 and 1.8, convergence in the scaling limit of the UST on $\mathbb{Z}^{3}$ now holds for an arbitrary scaling. This convergence is established with respect to Gromov-Hausdorff-Prohorov topology, incorporating a locally uniform topology for the embedding (for a detailed presentation of this topology, refer to [1]).
@@ 1.2 Proof overview
We now comment on the one-point estimates. Recall that in [19], for obtaining the strong one-point function asymptotics, the authors compared the non-intersection probability of independent LERW and SRW started at the origin, with the non-intersection probability of independent LERW and SRW with separated starting points; the gap between these starting points allowed the comparison of the LERW with its scaling limit and the non-intersection probability with respect to an independent Brownian motion. To extend the one-point estimate (i.e., Theorem 1.3) beyond dyadic scales, we examine scales of the form $m=2^{n+r}$ with $n\in\mathbb{N}_{+}$ and $r\in[0,1)$. Given a fixed $x\in\mathbb{D}\setminus\{0\}$, we consider a function $f:[0,1]\rightarrow\mathbb{R}^{+}$ satisfying $P(x_{2^{m}}\in\eta_{2^{m}})\simeq f(r)2^{-(3-\beta)n}$. In fact, the function $f$ satisfies a certain functional equation (stated in Proposition 4.1) which implies that $f(r)=c2^{-(3-\beta)r}$ for some $c>0$. Therefore $P(x_{2^{m}}\in\eta_{2^{m}})\simeq c2^{-(3-\beta)m}$ for any $m\in\mathbb{R}^{+}$ and yields Theorem 1.3. The key part property of the functional equation is a local logarithmic property, which is proved in Proposition (4.2) with the techniques in [19] mentioned above.

### v2

@@ 1 Introduction
In considering the scaling limit of discrete random curves, it is natural to rescale the space by the inverse of the lattice mesh, parametrize the discrete curve in a way such that each lattice edge is traversed in the same amount of time and rescale the time in a way such that a smooth bounded domain is traversed in time of order 1. If the limit of the rescaled and parametrized random curve exists, we refer to it as the scaling limit in natural parametrization. Recently, the last two authors proved in [20, 22] that when rescaled properly, three-dimensional LERW converges as a parametrized curve to its scaling limit in natural parametrization and determined in [21] the optimal Hölder exponent of the scaling limit.
With these results in mind, it is natural to ask if there is a more direct and canonical way to obtain the natural parametrization of the scaling limit. In the 2D case, Lawler and Viklund showed in [18, 19] that the scaling limit of planar LERW in natural parametrization is given through the Minkowski content of ${\rm SLE}_{2}$, whose existence was confirmed in [15]. It has also been shown (see [17, 3]) that the natural parametrization of ${\rm SLE}_{k}$ (which satisfies a conformal Markov property) is unique up to multiplicative constants. However, conceivably it is difficult to establish such axiomatic characterization of natural parametrization in 3D due to lack of the rich conformal symmetry in 2D.
In this work we are going to show that the Minkowski content of the scaling limit of 3D LERW exists and is equivalent to the limiting occupation measure obtained by the authors in [22] from which the natural parametrization of the scaling limit is induced. In the course of proof, we also obtain sharp asymptotics of one-point function in any scale, and sharp estimates on the ball-hitting probabilities of LERW and its scaling limit, which is a big improvement of earlier results of the authors in [20].
@@ 1.1 Main results
We now turn to the main results of this paper. Before stating them precisely, we need to introduce some notation first. Let $\mathbb{D}=\{x\in\mathbb{R}^{3}\ |\ |x|<1\}$ be the open unit ball. Pick some $m\in\mathbb{R}^{+}$, and let $\mathbb{D}_{m}\coloneqq m^{-1}\mathbb{Z}^{3}\cap\mathbb{D}$ be the discretized unit ball on the rescaled integer lattice $m^{-1}\mathbb{Z}^{3}$.
Let $S_{m}$ be the simple random walk on $m^{-1}\mathbb{Z}^{3}$ started from 0 and stopped at the first exit of $\mathbb{D}_{m}$. Define $\eta_{m}\coloneqq{\rm LE}(S_{m})$ as its loop-erasure, which we will refer to as the LERW on $\mathbb{D}_{m}$. (We define the loop-erasure of a path below, in Definition 2.1.) Write ${\cal K}_{m}$ for the trace of $\eta_{m}$ as a random compact subset of $\mathbb{R}^{3}$, and let $\cal K$} stand for the weak limit of ${\cal K}_{2^{n}}$ as $n\in{\mathbb{Z}^{+}}$ tends to infinity, whose existence was confirmed in [11]. Let
@@
be the reference scaling factor. Here, $\hat{x}_{m}$ stands for one of the nearest points to $\hat{x}$ among $m^{-1}\mathbb{Z}^{3}$. From up-to-constant one-point estimates, we have $f_{m}\asymp m^{\beta}$ (see Subsection 2.6 below). Let
$\mu_{m}\coloneqq(f_{m})^{-1}\sum_{x\in\mathbb{D}_{m}\cap\eta_{m}}\delta_{x}$ (1.3)
@@
It was proved in [22, Theorem 1.1] that as $m\in{\mathbb{R}^{+}}$ tends to infinity,
@@
for some random measure $\mu$ supported on $\cal K$}, which is measurable with respect to $\cal K$}. We call $\mu$ the limiting occupation measure.
@@
From now on, we regard the LERW $\eta_{m}$ as a parametrized curve where each edge of $m^{-1}\mathbb{Z}^{3}$ is traversed in time $(f_{m})^{-1}$ at a constant speed. Let $\eta$ be the random continuous curve obtained by parametrizing $\cal K$} with the limiting occupation measure $\mu$, as described in [22, Subsection 9.1]. It was also proved in [22, Theorem 1.3] that $\eta_{m}$ converges to $\eta$ in natural parametrization. In other words, as $m\in{\mathbb{R}^{+}}$ tends to infinity,
@@
We now turn to Minkowski content. Roughly speaking, the Minkowski content of a set is defined through the limit of the renormalized volume of the enlargement of the set. Although not a measure per se, it is a natural way to define “fractal” measures on fractal sets. Given a Borel set $G\subset{\mathbb{R}^{3}}$ and $\delta\in(0,3]$, we write ${\rm Cont}_{\delta}(G)$ for its $\delta$-dimensional Minkowski content (provided it exists); see Subsection 2.3 for its precise definition. Let ${\cal D}^{o}$ be the collection of “nice” dyadic boxes defined in (2.12).
@@
There exists a universal constant $c_{0}>0$ such that for any dyadic box $V\in{\cal D}^{o}$, ${\rm Cont}_{\beta}({\cal K}\cap V)$ exists and equals to $c_{0}\mu(V)$ almost surely.
By a standard measure-theoretic procedure, see Proposition 2.2 below, we can generate a random non-atomic Borel measure $\nu$ out of the Minkowski content of $\cal K$}.
@@
The Borel measure $\nu$ is supported on $\cal K$} and defines an intrinsic parametrization. Given $x\in{\cal K}$, denote by ${\cal K}_{x}\subset{\cal K}$ the trace of the curve between 0 and $x$. For each $t\in[0,\nu({\cal K})/c_{0}]$, there uniquely exists $x_{t}\in{\cal K}$ such that $t=\nu({\cal K}_{x_{t}})/c_{0}$, and then we define
@@
The relevance of Corollary 1.2 is that the parametrization of $\eta$, which was obtained via the scaling limit of the occupation measure $\mu$ (which is a priori not necessarily scaling-invariant), is equal to an intrinsic parametrization of $\cal K$}, up to a universal constant factor. Moreover, since the measure $\nu$ is both scaling-invariant and rotation-invariant, then the parametrized limit curve $\eta$ is scaling-invariant and rotation-invariant as well (see Corollary 6.5).
In the course of the proof of Theorem 1.1, we also obtain accurate asymptotics for any scale of the probability that LERW passes a certain point or a certain ball which we refer to loosely as “one-point functions”. We present these results here, since both of them are of independent interest.
@@
We now turn to the continuum and look at the probability of intersecting a ball. For $x\in\mathbb{R}^{3}$ and $r>0$, let $B(x,r)\coloneqq\{y\in\mathbb{R}^{3}\,|\,|x-y|\leq r\}$. Elaborating on Theorem 1.3 and various coupling techniques developed so far, we are able to show the following asymptotics of the probability that the scaling limit $\cal K$} hits a ball.
@@
There exist $c_{1},c,\delta>0$ such that for any $x\in\mathbb{D}\setminus\{0\}$, for $r<\min\{|x|,1-|x|\}/2$,
$P\Big({\cal K}\cap B(x,r)\neq\emptyset\Big)=c_{1}g(x)r^{3-\beta}\Big[1+O\big(d_{x}^{-c}r^{\delta}\big)\Big],$ (1.9)
@@
Note that $g(x)$ here corresponds to $c(x)$ in [20, Theorem 1.1] (which we rephrase in Proposition 2.11), where only dyadic scales were considered. Hence, $g(x)$ satisfies the following asymptotics (see [20, Equations (1.5), (1.6)]):
@@
Moreover, for any $(z,w)\in\Omega$ and $r<d_{z,w}/2$
$P\left({\cal K}\cap B(z,r)\neq\emptyset,{\cal K}\cap B(w,r)\neq\emptyset\right)=c_{1}^{2}g(z,w)r^{2(3-\beta)}\Big[1+O(d_{z,w}^{-c}r^{\delta})\Big],$ (1.15)
@@
Theorem 1.1 and Theorem 1.3 enhance [22, Theorem 1.3]. Specifically, the refined one-point estimates allow us to substitute the reference scaling factor (1.2) with an explicit scaling. We define
$\overline{\mu}_{m}\coloneqq m^{-\beta}\sum_{x\in\mathbb{D}_{m}\cap\eta_{m}}\delta_{x}$ (1.16)
and introduce $\overline{\eta}_{m}$ as the LERW on $\mathbb{D}_{m}$ parametrized by $\overline{\mu}_{m}$ and $\overline{\eta}$ as $\cal K$} parametrized by $(c_{0}g(\widehat{x}))^{-1}$-times its Minkowski content. We now present the convergence of 3D LERW in its natural parametrization in full strength.
@@
As $m\to\infty$, $\overline{\eta}_{m}\xrightarrow{\rm w}\overline{\eta}$ in the topology generated by the $\rho$-metric.
@@
Let $\mathcal{C}$ be the space of parametrized curves $\lambda:[0,\infty)\to\mathbb{R}^{3}$ satisfying $\lambda(0)=0$ and $\lim_{t\to\infty}{\rm diam}(\lambda[0,t])=\infty$. Consider the metric on $\mathcal{C}$ given by
@@
Let $S_{m}$ be a simple random walk on $m^{-1}\mathbb{Z}^{3}$ started from 0. The path
@@
defines the infinite loop-erased random walk (ILERW) on $m^{-1}\mathbb{Z}^{3}$. Since the simple random walk on $\mathbb{Z}^{3}$ is transient almost surely, then $\eta^{\infty}_{m}$ is well-defined. We also write $\eta^{\infty}_{m}$ for the curve obtained by linearly interpolating the path $\eta^{\infty}_{m}$, and consider a parametrization $\overline{\eta}^{\infty}_{m}$ defined by
@@
1) Using the same techniques employed in proving Theorem 1.3, one can show the existence of the one-point function associated with $\eta^{\infty}$, the scaling limit of the ILERW. Specifically, letting $g_{\infty}:=b_{1}|x|^{-(3-\beta)}$ (recall the definition of $b_{1}$ from Corollary 1.5), using an asymptotic relation between LERW and ILERW (see e.g. Lemma 3.1 of [22]), there exist $c,\delta>0$ such that for any $x\in\mathbb{D}\setminus\{0\}$,
$P\Big(x_{m}\in\eta_{m}^{\infty}\Big)=g_{\infty}(x)m^{\beta-3}\Big[1+O\Big(|x|^{-c}m^{-\delta}\Big)\Big]\qquad(\text{as }m\to\infty).$ (1.17)
Moreover, for all $r\in(0,|x|/2)$,
$P\left(\eta^{\infty}\cap B(x,r)\neq\emptyset\right)=c_{1}g_{\infty}(x)r^{3-\beta}\Big[1+o_{r}(1)\Big],$
where $c_{1}$ is defined in Theorem 1.4. In a similar fashion, we can also define the two-point functions $g_{\infty}(\cdot,\cdot)$. We omit the details.
2) Although open at the moment, we conjecture that $g(\cdot)$ is differentiable and $|g^{\prime}(\cdot)|$ satisfies certain asymptotics as $x\to 0$ or $x\to 1$.
3) Under certain regimes, by performing an analysis similar to those in this work, it is possible to relate the asymptotics of $g(\cdot,\cdot)$ and $g_{\infty}(\cdot,\cdot)$ to those of $g(\cdot)$ or $g_{\infty}(\cdot)$. For example, as $x\to 0$ or $x\to 1$, for any $y\in\mathbb{D}\setminus\{0\}$, one has
$\frac{g(x,y)}{g(x)}\to g(y).$ (1.18)
This is reminiscent of the operator product expansion (OPE) in 2D conformal field theory.
@@
1) The existence of a scaling limit for the uniform spanning tree on ${\mathbb{Z}}^{3}$ was established in [1]. The paper relies on the scaling limit of the loop-erased random walk in [22] and the one-point function estimates for the loop-erased random walk in [20]. Building upon the results from these papers, [1] shows the existence of sub-sequential scaling limits, convergence over a dyadic scaling, and properties of the limit tree.
As noted in [1, Remark 1.2], the restrictions of [22] and [20] to a dyadic scaling impose the same condition over the results in [1]. With the introduction of Theorems 1.31.7, and 1.8, the convergence of the scaling limit of the UST on $\mathbb{Z}^{3}$ now holds for arbitrary scaling. This convergence is established with respect to the Gromov-Hausdorff-Prohorov topology, incorporating a locally uniform topology for the embedding. For a detailed presentation of this topology, see [1].
@@ 1.2 Proof overview
We now comment on the one-point estimates. Recall that in [20], for obtaining the strong one-point function asymptotics, the authors compared the non-intersection probability of independent LERW and SRW started at the origin, with the non-intersection probability of independent LERW and SRW with separated starting points; the gap between these starting points allowed the comparison of the LERW with its scaling limit and the non-intersection probability with respect to an independent Brownian motion. To extend the one-point estimate (i.e., Theorem 1.3) beyond dyadic scales, we examine scales of the form $m=2^{n+r}$ with $n\in\mathbb{N}_{+}$ and $r\in[0,1)$. Given a fixed $x\in\mathbb{D}\setminus\{0\}$, we consider a function $f:[0,1]\rightarrow\mathbb{R}^{+}$ satisfying $P(x_{2^{m}}\in\eta_{2^{m}})\simeq f(r)2^{-(3-\beta)n}$. In fact, the function $f$ satisfies a certain functional equation (stated in Proposition 4.1) which implies that $f(r)=c2^{-(3-\beta)r}$ for some $c>0$. Therefore $P(x_{2^{m}}\in\eta_{2^{m}})\simeq c2^{-(3-\beta)m}$ for any $m\in\mathbb{R}^{+}$ and yields Theorem 1.3. The key part property of the functional equation is a local logarithmic additivity property, which is proved in Proposition 4.2 with the techniques in [20] mentioned above.
1.3 Organization of this work